The difference between the sides containing right angle in the right angles triangle is 7 cm. The ares of the triangle is 60 cm2 . Find its perimeter

Grade:

4 Answers

shivansh
9 years ago
let one side b x. then other side wil be x+7. now 1/2(x+7)(x)=60. by solving d quadratic eqution , u wil get x=8 or x= -15. as sides cnt be -ve, so x=8. nw one side is 8 nd another is 8+7=15. nw using pythagoras theorm, hypotaneous=17. so perimeter=15+8+17=40 ans
Akshay Sreekumar
9 years ago
40
Gursimran Singh
9 years ago
firstly let the base= a and the altitude= a+7(or a-7) now using the formula area of triangle=1\2*base*height solve this equation in as
1\2(a)(a+7)=60
a^2+a7\2=60
a^2+a7=120
a*(a+7)=120
a=120\a+7
and on futher calculation you get a=8 by splitting it again and finding value of a so base =8 and height or altitude=8+7=15
1\2*15*8
=15*4
=60
hence proved
Gursimran Singh
9 years ago
and after it you can find perimeter firstly let x= hypotenuse by making equation x^2=under root 15^2 +8^2
=under root 225+64=under root 289 and we get the hypotenuse 17 now add these 17+8+15=40 ans actually i forgot to give answer in earlier post sorrry